OpenStudy (anonymous):

What is the x-intercept for this linear equation? 6x-y=9

OpenStudy (luigi0210):

The x intercept is when y=0

OpenStudy (luigi0210):

So set your y equal to zero and solve

OpenStudy (anonymous):

but they want the x-intercept

OpenStudy (luigi0210):

6x-y=9 6x-0=9 6x=9 x=3/2

OpenStudy (anonymous):

and how do i wirte that in a decimal

OpenStudy (luigi0210):

1.5

OpenStudy (anonymous):

k